Your company sponsors a 401(k) plan into which you deposit 12 percent of your $75,000 annual income. Your company matches 50 percent of the first 3 percent of your earnings. You expect the fund to yield 8 percent next year. If you are currently in the 31 percent tax bracket. a. How many dollars did you invest out of your salary in your 401(k) plan this year? Annual investment $ b. What is your one-year return? (Round your answer to 2 decimal places. (e.g., 32.16)) One-year return %
Answer #1

a: Amount invested out of salary = 12%*75000 = 9000

b: Income = (Amount invested by self + amount invested by employer) * Yield

= (9000+ 50%*3%*75000) * 8%

= 810

One year return before tax= Income/ Investment

= 810/9000

= 9%

One year return post tax = 9%*(1-0.31) = 6.21%